OpenStudy (anonymous):

Please help 3 easy questions! Solve for x: 5x + 2 = 4x − 9 7 −7 11 −11 Solve for x: 3(x + 1) = −2(x − 1) − 4 1 −1 −5 −25

OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (anonymous):

ok ok i will help

OpenStudy (anonymous):

ok so how do i solve these?

OpenStudy (anonymous):

ok so for number one (i will explain them) \[5x +2 = 4x -9\] \[9x +2 =-9\] then you minus two from both sides and they you will come out with \[9x= -7\]

OpenStudy (anonymous):

do you know how to solve them??????

OpenStudy (anonymous):

it

OpenStudy (anonymous):

ok well i will show you \[x = -1.28\]

OpenStudy (anonymous):

o wait that is not there